OpenStudy (anonymous):

Please help me I will give medal and become fan if you explain. Find the equation of the graph that is parallel and perpendicular that passes through the specified point. 4x + 3y = 12 (-1,-1)

OpenStudy (anonymous):

@esshotwired Do you know how to do this?

OpenStudy (anonymous):

@kewlgeek555 @lauluisa @timaashorty @InYourHead @brendalopez @e.cociuba @TuringTest

OpenStudy (esshotwired):

so u need 2 lines that pass through (-1, -1)?? parallel and perpendicular?

OpenStudy (anonymous):

I guess lol I am so confused and that was all that the question gave me.

OpenStudy (esshotwired):

umm ok well...im gunna think to myself and u just read it until i ask u a question

OpenStudy (anonymous):

k

OpenStudy (esshotwired):

4x + 3y = 12 3y=-4x+12 y=-4/3x+12 parallel: -4/3x perp: 3/4x

OpenStudy (esshotwired):

y+1=-4/3(x+1) y+1=-4/3x-4/3 y=-4/3x-7/3 parallel ^^

OpenStudy (esshotwired):

y+1=3/4(x+1) y+1=3/4x+3/4 y=3/4x+7/4 perpendicular ^^

OpenStudy (anonymous):

how you get that? The last one?

OpenStudy (the_fizicx99):

@esshotwired perp is the reciprocal and negative i.e 3x -> -1/3x

OpenStudy (esshotwired):

the slope of a line perpendicular to another line is the opposite reciprocal. meaning you flip the fraction and times it by -1. if it is not a fraction then the number will be the denominator with a 1 on top.

OpenStudy (anonymous):

oh but where did the (x=1) come from?

OpenStudy (esshotwired):

you mean (x+1)? it is part of the point slope form: \[y-y ^{1}=slope(x-x ^{1})\]